OpenStudy (anonymous):

Eliminate parameter t, x = 5 sin t, y = 2 sin t. Do I need to find cos t for one of the functions?

OpenStudy (anonymous):

t = arcsin(x/5) y = 2 sin(arcsin(x/5)) y = 2 * x / 5

OpenStudy (anonymous):

sorry. but that doesnt make any sense to me. wasn't looking for answer, but more a path in the right direction.

ganeshie8 (ganeshie8):

solve "sint" from first equation, then plug it in the second equation

ganeshie8 (ganeshie8):

you have : x = 5sint sint = ?

OpenStudy (anonymous):

got it. thanks

ganeshie8 (ganeshie8):

yw